OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle if a = 3 and b = 4. Use: a2 + b2 = c2 5 Square root of 7 25 7

OpenStudy (johnweldon1993):

Right, so you have 3 sides of the triangle.... a = one side b = another side c = the hypotenuse (longest side) if you have 2 sides of the triangle and want to solve for the other...you can use that equation a² + b² = c² so you can plug in 3 for 'a' 4 for 'b' and solve for c 3² + 4² = c² so to solve for c \[c = \sqrt{3^2 + 4^2}\]

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

√25 now

OpenStudy (johnweldon1993):

what you had so far was 3² + 4² ...that does = 25 now you do the square root of that 25....

OpenStudy (anonymous):

Thanks.... Could you help me with a couple more? @johnweldon1993

OpenStudy (johnweldon1993):

if I can then yes :)

OpenStudy (anonymous):

Would the square root of 5 be 5?

OpenStudy (johnweldon1993):

the square root of 25 is 5... the square root of 5 is NOT 5

OpenStudy (anonymous):

Your confusing me...

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

This is still on that first question correct?

OpenStudy (anonymous):

Yeah... So the answer is 5 not 25?

OpenStudy (johnweldon1993):

The answer is 5 yes

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

\[c = \sqrt{3^2 + 4^2}\] \[c = \sqrt{9 + 16}\] \[c = \sqrt{25}\] \[\sqrt{25} = 5\] better?